Directrix (directrix):

A Calculus person is needed here: What am I doing when I take a triple integral? I know a double integral of a function is finding volume?

Directrix (directrix):

http://openstudy.com/study#/updates/52954d2ee4b0f4040dd3e824 See second question at bottom of thread above. Thanks.

OpenStudy (anonymous):

triple integrals can also be finding volume. \[\int\limits \int\limits \int\limits dV = V\]

OpenStudy (anonymous):

a single integral can also be used to find volume. A triple integral finds the area under a curve that is a function of three variables. The type of integral doesn't determine the unit, the units of the variables determine the final unit.